OpenStudy (anonymous):

@AlexandervonHumboldt2 anyone helpme again ?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

question?

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }-\frac{ 1 }{ 2}*\frac{ 7 }{ 8 }\] @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

1/2*7/8=(1*7)/(2*8)=????????

OpenStudy (anonymous):

2/16

OpenStudy (alexandervonhumboldt2):

no 7/16

OpenStudy (alexandervonhumboldt2):

3/4 - 7/16=12/16 - 7/16=??????

OpenStudy (anonymous):

oh sorry I put 1/2 sorry

OpenStudy (anonymous):

5/16

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

then what I do im sorry im bad in math im trying my best though

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@allylovesgumby @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

5/16 is anwer